OpenStudy (anonymous):

x^4-6x^3-7x^2=0 ,x1=?,x2=?,x3=?,x4=?

OpenStudy (anonymous):

So you factor this into: x^2(x^2-6x-7)=0 factor what is inside the parenthesis and get: x^2(x-7)(x+1)= The double root means you will have 0 twice, 7. and -1 as the roots

OpenStudy (anonymous):

what does that mean double root could u plz explain

OpenStudy (cwrw238):

the double root is 0 x^2 = 0